Grade 12th passMechanics

A unit vector is expressed as n cap=0.4i+0.8j+ck. The value of c must be.
Please explain.

Ishan karki
8 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear Ishan
 
Unit vector means its modulus is 1
hence
sqrt (0.16+0.64+c^2) = 1
0.80 + c^2 = 1
c^2 = 0.20
c = ± sqrt(0.20)
 
Regards
Arun (askIITians forum expert)